OpenStudy (anonymous):

for the parabola x^2+2x+28y-111=0. find the equation of the axis and what is the maximum value ?

OpenStudy (anonymous):

yeah thanks i got it :)!

OpenStudy (anonymous):

seperate x and y terms x^2+2x=111-28y complete square for x terms x^2+2x+1-1=111-28y (x+1)^2=112-28y writing in standard form (x+1)^2=-28(y-4) compare with standard equation (x-h)^h=4p(y-k) for this standard parabola x=h is it axis so x=-1 is the parabola axis and for maximium vale one method is to take derivative ( i assume you dont know it yet ) . other way is since this parabola is concave down so the max value is the y value of the vortex which is 4 so y=4 is the maximium hieght.

OpenStudy (anonymous):

welcome .:)